# Leibniz’s contested infinitesimals: Further depictions

### v1

@@ 2.1. Accidental and absolute impossibility
For Leibniz, fictional did not necessarily mean contradictory. Crucial here is Leibniz’s distinction between absolute impossibility and accidental impossibility. Already in the Confessio Philosophi, [^12] Leibniz [40], 1672/3, p. 128. Leibniz speaks of the concept of “impossible by accident” and contrasts it with “absolute impossibility” i.e., contradiction. He gives the examples of a species with an odd number of feet, and an immortal mindless creature, which are, according to him, harmoniae rerum adversa i.e., “contrary to the harmony of things,” [^13] Leibniz as translated by Sleigh in [48, p. 57]. but not necessarily contradictory.
@@
“a construction in our characters” (in nostris characteribus), [^14] Leibniz [41, p. 520].
meaning that one can carry out a formal calculation, such as those with imaginary roots, regardless of whether the mathematical notions involved idealize anything in nature. Leibniz goes on to discuss in detail the cases of imaginary roots and infinitesimals, which to him constitute examples of accidental impossibility. [^15] For a detailed textual analysis, see [33, Section 5.3]. The distinction was also analyzed by Eklund [18].
@@ 2.2. An epigraph and an artful ellipsis
Accordingly, genuine infinitesimals are fine in mathematics just as imaginary roots are, but insisting on their existence beyond the realm of mathematics would push them beyond common sense, according to Leibniz. By suppressing the crucial middle sentence, AR changed the apparent meaning of the Leibnizian passage. The letter to Fontenelle proves precisely the opposite of what AR make it out to say by means of a skillful use of an ellipsis. [^23] Earlier skillful uses of the truncation technique to score an apparently syncategorematic point in the writings of Ferraro [22, p. 29], Goldenbaum [24, p. 76], Ishiguro [28], and Tho [61, p. 71] were analyzed in Bascelli et al. [11].
@@ 3. Infinite wholes and bounded infinities
[Leibniz] distinguished between two infinites, the bounded infinite straight line, the recta infinita terminata, and the unbounded infinite straight line, the recta infinita interminata. He investigated this distinction in several studies from the year 1676. Only the first kind of straight lines can be used in mathematics, as he underlined in his proof of theorem 11 [i.e., Propositio XI]. [^26] Knobloch [36, p. 97].
Leibniz mentioned the distinction in a 29 july 1698 letter to Bernoulli. Here Leibniz analyzes a geometric problem involving unbounded infinite areas and states an apparent paradox: “Therefore the two infinite spaces are equal, and the part is equal to the whole, which is impossible.”272727“Ergo haec duo spatia infinita sibi sunt aequalia, pars toti, quod est absurdum” (Leibniz [45, p. 523]). To resolve the paradox (i.e., the clash with the part-whole principle), Leibniz exploits the notion of bounded infinity. Far from creating paradox, bounded infinities serve to resolve one. [^28] For a detailed textual analysis, see [35, Section 1]. Further details on infinita terminata appear in Section 7.
@@ 3.3. Conservativity results by Friedman and others
In the late 1960s, Harvey Friedman proved the conservativity of not only PA over PA, but the more powerful theory, namely PA plus the induction scheme referred to as ${}^{*}\Pi_{\infty}\mathsf{-Induction}$. Friedman’s work was mentioned by Kreisel. [^35] See Kreisel [38, p. 94]. See also Feferman [21, p. 967] (second theorem in Section 8.7) and Kossak [37]. Friedman’s result was followed up by Keisler et al. [26]. Thus, Robinson’s question concerning axiomatisations was answered by Friedman and led to a mature body of work. [^36] See Enayat [19] for further details.
@@
[One] notes that Leibniz stresses the distinction between infinite cardinality and infinite quantity (reciprocal of infinitesimals) [and] argues that the exchange with Bernoulli precisely refutes [the] attempt to blend infinite cardinality and infinite quantity so as to deduce the inconsistency of infinitesimals. [^37] Katz et al. [34, p. 264].
@@
However, the authors of [34] claimed no such thing. The distinction is not between discrete and continuous quantity or magnitude, contrary to what is claimed by AR. The distinction is not between number and magnitude either, contrary to what they claim. Rather, the distinction is the Leibnizian distinction between magnitude and multitude; [^40] For details see e.g., [33, Section 4.1]. in other words, between magnitude or quantity (as formalized by unlimited numbers in modern theories e.g., PA, as discussed in Section 3.3), on the one hand, and infinite wholes formalized as Cantorian infinities (whether ordinal or cardinal), on the other. [^41] Leibniz has a more general notion of an ‘infinite whole’ that includes unbounded lines and unbounded regions. See the beginning of Section 3 and sources given there for a discussion of the paradox of unbounded regions in an exchange of letters with Bernoulli. If there is any invention involved, it is due to AR rather than the authors of [34]. Such a misunderstanding and an attendant conflation are a persistent feature of AR’s writing on Leibniz, as when they write:
@@ 4. Homogeneity and comparability
Note that this Leibnizian definition mentions both number and fraction. By AR’s logic, fractions would similarly be incompatible with infinitesimals. If so, how would AR account for Leibniz’s definition of infinitesimal as a fraction, namely “[an] infinitely small fraction, or one whose denominator is an infinite number”474747“fraction infiniment petite, ou dont le denominateur soit un nombre infini” Leibniz [47, p. 93]. ?
@@
Regrettably, AR keep ignoring the fact that in 1695 Leibniz made it clear in two separate texts that infinitesimals violate Euclid’s Definition V.4.484848Leibniz [43, p. 288], [44, p. 322]. For a detailed analysis see [8, Section 1.5, pp. 177–179]. Neither of these Leibnizian texts is even cited in [6], even though AR appear to acknowledge that the presence of inassignables would signal non-Archimedean behavior; see Section 5.4. Instead, they offer non-sequiturs, such as the following:
@@ 4.2. Ishiguro on referring and designating
The parenthetical reference is to Leibniz [47, p. 93]. In her Chapter 5, Ishiguro seeks to create the impression that the letter to Varignon provides evidence for her interpretation. Does Leibniz apply the term syncategorematic to infinite magnitudes in this letter? We will examine the issue in Section 4.5.
@@ 4.4. Arthur’s primary sources
There is a brief “supplementary study” [49] where Leibniz seems to illustrate syncategorematic actuality by numbers, but it turns out that the study was “Written on a separate slip of paper and crossed through by Leibniz” [53, p. 409]; furthermore, the supplementary study was never sent to des Bosses. It seems ill-advised to use such a flimsy study as a basis so as to re-evaluate (as Arthur attempts to do in [5, Section 10.5]) Leibniz’s well-established position concerning actuality as applicable only to actual things.
@@ 4.5. Analysis of letter to Varignon
Yet we must not imagine that this explanation debases the science of the infinite and reduces it to fictions, for there always remains a ‘syncategorematic’ infinite, as the Scholastics say … [^66] Leibniz [47, p. 93] as translated by Loemker (p. 543).
@@ 5.2. Transfer principle
Leibniz did say, in one of the passages quoted above, that what succeeds for the finite numbers succeeds also for the infinite numbers and vice versa, and this is remarkably close to our transfer of statements from ${\mathbb{R}}$ to ${{}^{\ast}{\mathbb{R}}}$ and in the opposite direction. [^74] Robinson [59, p. 266].
@@ 5.4. Leibniz, Nelson, and inassignables
This phase in Leibniz’s development was analyzed in detail by Ugaglia and Katz. [^79] Ugaglia and Katz [62, Section 3.2.1]. During this period,
[E]xtension is continuous, and continuity means for Leibniz potentially infinite divisibility, that is syncategorematicity, as it was for Aristotle. The terminology itself of unassignability and infinitesimals is absent in this phase: [^80] Two years later, Leibniz will use inassignabilis in his Theoria motus abstracti [39, p. 265]. This text belongs to Phase II of Leibniz’s development. ‘infinitely small’ just means ‘potentially infinitely divisible.’ [^81] Ugaglia and Katz [62, p. 13].
@@ 6. Hornangles vs inassignables
[A]lready Eudoxus knew that also other number systems – so-called non-Archimedean number systems – were conceivable … Such a system of quantities which was already known to the Greeks were cornicular angles or horn angles. [^84] Sonar [60, p. 40].
Surprisingly, AR open their section 7 on hornangles with yet another conflation of non-Archimedean phenomena and ‘nonstandard quantities’, as follows: “it might be useful to analyse a last example regularly mentioned in support of the existence of nonstandard quantities in Leibniz: horn angles.” [^85] Arthur and Rabouin [6, p. 39]; emphasis added. AR fail to provide any quotation of an alleged attribution of a “nonstandard” attribute to hornangles, and with good reason: we are not aware of any such attribution. [^86] The reference [33, Section 4.5] cited by AR in footnote 99 on page 39 of [6] does contain a discussion of hornangles, but certainly makes no such attribution, contrary to AR’s claim.
@@ 6.1. Euclidis Prôta: page 191
AR proceed to contest the conclusions of the analysis of Leibniz’s use of hornangles as developed by Katz et al. [33, Section 4.5]. AR rely mainly on these two quotations from page 191 in GM to develop their rebuttal. While they mention the year, 1712, of the Leibnizian text Euclidis Prôta, they omit to mention the year of the Leibnizian text that provided the basis for the analysis in [33]. That year is 1686 (Leibniz [42]), namely no fewer than 26 years prior to Euclidis Prôta.
@@ 6.2. Leibniz on Peletier and Clavius
“And so in this matter I agree with Peletarius against Clavius; and Euclid, when he said that the angle of contact is smaller than any straight line, spoke a little too loosely, understanding “smaller” to mean something whose beginnings fall within the space of the former. But it should not therefore be considered that he has attributed a perfect quantity to the angle of contact with respect to the straight line. … and Clavius erred greatly, when he denied this axiom, by which it is asserted that what passes from one extreme to another, must also equally pass through everything in between, and on this account gave Thomas Hobbes an opportunity to insult geometers.” [^90] Leibniz [50, p. 191]; translation ours.
@@ 9. Two syncategorematists against Bos
This first basic example was not mentioned in Henk Bos’s seminal study (Bos 1974–5), and this had the unfortunate consequence of hiding the global structure of the reasoning.103103103Rabouin and Arthur [58, p. 439]; emphasis added.
@@ 10. Arthur vs Rabouin: Limit processes in 2013
A profound evolution of the concept involved in differential calculus, already emphasized by D’Alembert, was the progressive introduction of a new interpretation of its basic constituents in terms of “limit processes” – an interpretation that is absent from the practice of Leibniz and his first followers, but present in the Newtonian tradition.115115115Rabouin [56, p. 24]; emphasis added. By 2017, the passage was modified by deleting the phrase “absent from the practice of Leibniz” as follows: “A profound evolution of the concept involved in differential calculus, already emphasized by Jean le Rond d’Alembert, was the progressive introduction of a new interpretation of its basic constituents in terms of limit processes–an interpretation that is put forward in the Newtonian tradition. The Leibnizian calculus developed quite rapidly as a mixture of features derived from Leibniz’s and Newton’s techniques, etc.” [57, p. 213]. The 2013 version was the more accurate one.

### v2

@@ 2.1. Accidental and absolute impossibility
For Leibniz, fictional did not necessarily mean contradictory. Crucial here is Leibniz’s distinction between absolute impossibility and accidental impossibility. Already in the Confessio Philosophi, [^12] Leibniz [42], 1672/3, p. 128. Leibniz speaks of the concept of “impossible by accident” and contrasts it with “absolute impossibility” i.e., contradiction. He gives the examples of a species with an odd number of feet, and an immortal mindless creature, which are, according to him, harmoniae rerum adversa i.e., “contrary to the harmony of things,” [^13] Leibniz as translated by Sleigh in [50, p. 57]. but not necessarily contradictory.
@@
“a construction in our characters” (in nostris characteribus), [^14] Leibniz [43, p. 520].
meaning that one can carry out a formal calculation, such as those with imaginary roots, regardless of whether the mathematical notions involved idealize anything in nature. Leibniz goes on to discuss in detail the cases of imaginary roots and infinitesimals, which to him constitute examples of accidental impossibility. [^15] For a detailed textual analysis, see [34, Section 5.3]. The distinction was also analyzed by Eklund [18].
@@ 2.2. An epigraph and an artful ellipsis
Accordingly, genuine infinitesimals are fine in mathematics just as imaginary roots are, but insisting on their existence beyond the realm of mathematics would push them beyond common sense, according to Leibniz. By suppressing the crucial middle sentence, AR changed the apparent meaning of the Leibnizian passage. The letter to Fontenelle proves precisely the opposite of what AR make it out to say by means of a skillful use of an ellipsis. [^23] Earlier skillful uses of the truncation technique to score an apparently syncategorematic point in the writings of Ferraro [22, p. 29], Goldenbaum [24, p. 76], Ishiguro [28], and Tho [63, p. 71] were analyzed in Bascelli et al. [11].
@@ 3. Infinite wholes and bounded infinities
[Leibniz] distinguished between two infinites, the bounded infinite straight line, the recta infinita terminata, and the unbounded infinite straight line, the recta infinita interminata. He investigated this distinction in several studies from the year 1676. Only the first kind of straight lines can be used in mathematics, as he underlined in his proof of theorem 11 [i.e., Propositio XI]. [^26] Knobloch [38, p. 97].
Leibniz mentioned the distinction in a 29 july 1698 letter to Bernoulli. Here Leibniz analyzes a geometric problem involving unbounded infinite areas and states an apparent paradox: “Therefore the two infinite spaces are equal, and the part is equal to the whole, which is impossible.”272727“Ergo haec duo spatia infinita sibi sunt aequalia, pars toti, quod est absurdum” (Leibniz [47, p. 523]). To resolve the paradox (i.e., the clash with the part-whole principle), Leibniz exploits the notion of bounded infinity. Far from creating paradox, bounded infinities serve to resolve one. [^28] For a detailed textual analysis, see [37, Section 1]. Further details on infinita terminata appear in Section 7.
@@ 3.3. Conservativity results by Friedman and others
In the late 1960s, Harvey Friedman proved the conservativity of not only PA over PA, but the more powerful theory, namely PA plus the induction scheme referred to as ${}^{*}\Pi_{\infty}\mathsf{-Induction}$. Friedman’s work was mentioned by Kreisel. [^35] See Kreisel [40, p. 94]. See also Feferman [21, p. 967] (second theorem in Section 8.7) and Kossak [39]. Friedman’s result was followed up by Keisler et al. [26]. Thus, Robinson’s question concerning axiomatisations was answered by Friedman and led to a mature body of work. [^36] See Enayat [19] for further details.
@@
[One] notes that Leibniz stresses the distinction between infinite cardinality and infinite quantity (reciprocal of infinitesimals) [and] argues that the exchange with Bernoulli precisely refutes [the] attempt to blend infinite cardinality and infinite quantity so as to deduce the inconsistency of infinitesimals. [^37] Katz et al. [35, p. 264].
@@
However, the authors of [35] claimed no such thing. The distinction is not between discrete and continuous quantity or magnitude, contrary to what is claimed by AR. The distinction is not between number and magnitude either, contrary to what they claim. Rather, the distinction is the Leibnizian distinction between magnitude and multitude; [^40] For details see e.g., [34, Section 4.1]. in other words, between magnitude or quantity (as formalized by unlimited numbers in modern theories e.g., PA, as discussed in Section 3.3), on the one hand, and infinite wholes formalized as Cantorian infinities (whether ordinal or cardinal), on the other. [^41] Leibniz has a more general notion of an ‘infinite whole’ that includes unbounded lines and unbounded regions. See the beginning of Section 3 and sources given there for a discussion of the paradox of unbounded regions in an exchange of letters with Bernoulli. If there is any invention involved, it is due to AR rather than the authors of [35]. Such a misunderstanding and an attendant conflation are a persistent feature of AR’s writing on Leibniz, as when they write:
@@ 4. Homogeneity and comparability
Note that this Leibnizian definition mentions both number and fraction. By AR’s logic, fractions would similarly be incompatible with infinitesimals. If so, how would AR account for Leibniz’s definition of infinitesimal as a fraction, namely “[an] infinitely small fraction, or one whose denominator is an infinite number”474747“fraction infiniment petite, ou dont le denominateur soit un nombre infini” Leibniz [49, p. 93]. ?
@@
Regrettably, AR keep ignoring the fact that in 1695 Leibniz made it clear in two separate texts that infinitesimals violate Euclid’s Definition V.4.484848Leibniz [45, p. 288], [46, p. 322]. For a detailed analysis see [8, Section 1.5, pp. 177–179]. Neither of these Leibnizian texts is even cited in [6], even though AR appear to acknowledge that the presence of inassignables would signal non-Archimedean behavior; see Section 5.4. Instead, they offer non-sequiturs, such as the following:
@@ 4.2. Ishiguro on referring and designating
The parenthetical reference is to Leibniz [49, p. 93]. In her Chapter 5, Ishiguro seeks to create the impression that the letter to Varignon provides evidence for her interpretation. Does Leibniz apply the term syncategorematic to infinite magnitudes in this letter? We will examine the issue in Section 4.5.
@@ 4.4. Arthur’s primary sources
There is a brief “supplementary study” [51] where Leibniz seems to illustrate syncategorematic actuality by numbers, but it turns out that the study was “Written on a separate slip of paper and crossed through by Leibniz” [55, p. 409]; furthermore, the supplementary study was never sent to des Bosses. It seems ill-advised to use such a flimsy study as a basis so as to re-evaluate (as Arthur attempts to do in [5, Section 10.5]) Leibniz’s well-established position concerning actuality as applicable only to actual things.
@@ 4.5. Analysis of letter to Varignon
Yet we must not imagine that this explanation debases the science of the infinite and reduces it to fictions, for there always remains a ‘syncategorematic’ infinite, as the Scholastics say … [^66] Leibniz [49, p. 93] as translated by Loemker (p. 543).
@@ 5.2. Transfer principle
Leibniz did say, in one of the passages quoted above, that what succeeds for the finite numbers succeeds also for the infinite numbers and vice versa, and this is remarkably close to our transfer of statements from ${\mathbb{R}}$ to ${{}^{\ast}{\mathbb{R}}}$ and in the opposite direction. [^74] Robinson [61, p. 266].
@@ 5.4. Leibniz, Nelson, and inassignables
This phase in Leibniz’s development was analyzed in detail by Ugaglia and Katz. [^79] Ugaglia and Katz [64, Section 3.2.1]. During this period,
[E]xtension is continuous, and continuity means for Leibniz potentially infinite divisibility, that is syncategorematicity, as it was for Aristotle. The terminology itself of unassignability and infinitesimals is absent in this phase: [^80] Two years later, Leibniz will use inassignabilis in his Theoria motus abstracti [41, p. 265]. This text belongs to Phase II of Leibniz’s development. ‘infinitely small’ just means ‘potentially infinitely divisible.’ [^81] Ugaglia and Katz [64, p. 13].
@@ 6. Hornangles vs inassignables
[A]lready Eudoxus knew that also other number systems – so-called non-Archimedean number systems – were conceivable … Such a system of quantities which was already known to the Greeks were cornicular angles or horn angles. [^84] Sonar [62, p. 40].
Surprisingly, AR open their section 7 on hornangles with yet another conflation of non-Archimedean phenomena and ‘nonstandard quantities’, as follows: “it might be useful to analyse a last example regularly mentioned in support of the existence of nonstandard quantities in Leibniz: horn angles.” [^85] Arthur and Rabouin [6, p. 39]; emphasis added. AR fail to provide any quotation of an alleged attribution of a “nonstandard” attribute to hornangles, and with good reason: we are not aware of any such attribution. [^86] The reference [34, Section 4.5] cited by AR in footnote 99 on page 39 of [6] does contain a discussion of hornangles, but certainly makes no such attribution, contrary to AR’s claim.
@@ 6.1. Euclidis Prôta: page 191
AR proceed to contest the conclusions of the analysis of Leibniz’s use of hornangles as developed by Katz et al. [34, Section 4.5]. AR rely mainly on these two quotations from page 191 in GM to develop their rebuttal. While they mention the year, 1712, of the Leibnizian text Euclidis Prôta, they omit to mention the year of the Leibnizian text that provided the basis for the analysis in [34]. That year is 1686 (Leibniz [44]), namely no fewer than 26 years prior to Euclidis Prôta.
@@ 6.2. Leibniz on Peletier and Clavius
“And so in this matter I agree with Peletarius against Clavius; and Euclid, when he said that the angle of contact is smaller than any straight line, spoke a little too loosely, understanding “smaller” to mean something whose beginnings fall within the space of the former. But it should not therefore be considered that he has attributed a perfect quantity to the angle of contact with respect to the straight line. … and Clavius erred greatly, when he denied this axiom, by which it is asserted that what passes from one extreme to another, must also equally pass through everything in between, and on this account gave Thomas Hobbes an opportunity to insult geometers.” [^90] Leibniz [52, p. 191]; translation ours.
@@ 9. Two syncategorematists against Bos
This first basic example was not mentioned in Henk Bos’s seminal study (Bos 1974–5), and this had the unfortunate consequence of hiding the global structure of the reasoning.103103103Rabouin and Arthur [60, p. 439]; emphasis added.
@@ 10. Arthur vs Rabouin: Limit processes in 2013
A profound evolution of the concept involved in differential calculus, already emphasized by D’Alembert, was the progressive introduction of a new interpretation of its basic constituents in terms of “limit processes” – an interpretation that is absent from the practice of Leibniz and his first followers, but present in the Newtonian tradition.115115115Rabouin [58, p. 24]; emphasis added. By 2017, the passage was modified by deleting the phrase “absent from the practice of Leibniz” as follows: “A profound evolution of the concept involved in differential calculus, already emphasized by Jean le Rond d’Alembert, was the progressive introduction of a new interpretation of its basic constituents in terms of limit processes–an interpretation that is put forward in the Newtonian tradition. The Leibnizian calculus developed quite rapidly as a mixture of features derived from Leibniz’s and Newton’s techniques, etc.” [59, p. 213]. The 2013 version was the more accurate one.